OpenStudy (anonymous):

y varies indirectly as twice x. When x = 2, y = 3. Find y when x = 6

OpenStudy (anonymous):

y=2kx where, k is constant plug , x =2 and y= 3 we have, 3 =2k(2) 3=4k 3/4 =k now we plug k =3/4 in y=2kx we have, y =2(3/4)x that is , y =3x/2

OpenStudy (anonymous):

for, x = 6 y =3(6)/2 =3(3) =9